OpenStudy (anonymous):

Will give out medals. Jeremiah’s scores on his math tests are 77, 80, 95, 91, and 77. What is the mean of Jeremiah’s scores?

OpenStudy (anonymous):

A. 77 B. 80 C. 84 D. 95

OpenStudy (anonymous):

84

OpenStudy (solomonzelman):

Mean is the average, add all terms. divide the result by the number of terms (in your case, by 5)

OpenStudy (anonymous):

Arithmetic Median Definition: Median is the middle value of the given numbers or distribution in their ascending order.Median is the average value of the two middle elements when the size of the distribution is even. Example 1: To find the median of 4,5,7,2,1 [ODD]. Step 1: Count the total numbers given. There are 5 elements or numbers in the distribution. Step 2: Arrange the numbers in ascending order. 1,2,4,5,7 Step 3: The total elements in the distribution (5) is odd. The middle position can be calculated using the formula. (n+1)/2 So the middle position is (5+1)/2 = 6/2 = 3 The number at 3rd position is = Median = 4 Example 2: To find the median of 4,5,7,2,1,8 [Even] Step 1: Count the total numbers given. There are 6 elements or numbers in the distribution. Step 2: Arrange the numbers in ascending order. 1,2,4,5,7,8 Step 3: The total elements in the distribution (6) is even. As the total is even, we have to take average of number at n/2 and (n/2)+1 So the position are n/2= 6/2 = 3 and 4 The number at 3rd and 4th position are 4,5 Step 4: Find the median. The average is (4+5)/2 = Median = 4.5

OpenStudy (solomonzelman):

Holy cra,... DUDE! he copy pasted it :) no one (even my dad) can't type this fast! LOL

OpenStudy (anonymous):

No my friend i am the fastest typer

OpenStudy (anonymous):

i just look it types

OpenStudy (solomonzelman):

Shut up dude, I saw how fast that was. A keyboard can't possible put up so many letters/symbols in 3 seconds :) Whatever dude

OpenStudy (anonymous):

hahahahaha just kidding

OpenStudy (solomonzelman):

Or perhaps you were typing and it didn't say you were? But even at that case, the Sherlock Holmes knows that you don't type 30 times faster than me although I do suck at typing :)

OpenStudy (anonymous):

i am faster at typeing but i might have big errors in my sentences